Transverse waves on a string have a wave speed 8.00 m/s, amplitude 0.0700 m, and


wavelength 0.320 m. The waves travel in the -x-direction, and at t = 0, x = 0, the end of


the string has its maximum upward displacement. (a) Find the frequency, period, and


wave number of these waves. (b) Write a wave function describing the wave. (c) Find


the transverse displacement of a particle at x = 0.360 m at time t = 0.150 s

1
Expert's answer
2022-01-21T17:06:33-0500

Given:

"v=8.00\:\rm m/s"

"A=\rm 0.0700\: m"

"\lambda=0.320\: \rm m"


(a) the frequency

"f=\frac{v}{\lambda}=\frac{8.00}{0.320}=25.0\:\rm Hz"

the period

"T=\frac{1}{f}=\frac{1}{25.0}=0.0400\:\rm s"

the wave number

"k=\frac{2\pi}{\lambda}=\frac{6.28}{0.320}=19.6\:\rm m^{-1}"

(b) the wave function describing the wave

"y(x,t)=A\cos(2\pi ft-kx)\\
y(x,t)=0.0700\cos (157t-19.6x)"

(c)


"y=0.0700\cos (157*0.150-19.6*0.360)\\
=-0.0495\:\rm m"